OpenStudy (anonymous):

Suppose that f(x) = sqrt(4 x - 7 ) g(x) = 8 x^2 - 5 . I need: h(x) = (g o g) h(x) = (f o f)

OpenStudy (jdoe0001):

h(x) = (gog) or (fof)?

OpenStudy (anonymous):

Yup g of g and f of f.

OpenStudy (anonymous):

I figured they'd be the same just replace them with themselves in their x's...like an inception. Seems I was mistaken.

OpenStudy (jdoe0001):

hmmm \(\bf f(x) = \sqrt{4x-7}\\ (fof)(x) = f(f(x)) = \sqrt{4(\sqrt{4x-7})-7}\) then

OpenStudy (jdoe0001):

\(\bf g(x) = 8 x^2 - 5\\ (gog)(x) = g(g(x)) = 8 (8 x^2 - 5)^2 - 5\)